Three resistors in a series connected to a battery

A 2.90 Ω resistor, a 5.79 Ω resistor, and a 9.96 Ω resistor are connected in series with a 15.0 V battery. What is the equivalent resistance?

What is the current in each resistor?
Answer #1

The three resistors are connected in series.

So, the equivalent resistance -

Req = R1 + R2 + R3 = 2.90 + 5.79 + 9.96 = 18.65 ohm.

In series combination, current through each resistor is the same.

And magnitude of current through each resistor = 15 / 18.65 = 0.80 A.